# Hierarchical Multi-agent Reinforcement Learning for Cyber Network Defense

### v2

@@ 5.3. H-MARL Transferability
In this section, we explore the possibility of knowledge transfer Weinshall et al. (2018); WEI et al. (2018); Zhuang et al. (2020); Nekoei et al. (2021) – given a defense strategy that has already been trained, can we use it to accelerate learning of a new defense that protects against a different adversary? Each of the sub-policies in our hierarchical architecture corresponds to a specific skill –recovering compromised hosts or investigating machines. We adapt the pre-trained sub-policy models to new attacks via fine-tuning, a well-established powerful transfer method for deep models Zhuang et al. (2020); Sun et al. (2019) that is significantly less costly than training from scratch. In our current design, we only fine-tune the sub-policies for a few iterations (between 5 and 10). Next, we fully train a new master policy, using the newly adapted sub-policies.
Figure 6 presents the fine-tuning results of the Investigate sub-policies, pre-trained against Aggressive Red, the average-performing attacker. A short fine-tuning is enough to adapt to a new adversary, resulting in learning curves similar to those obtained when training from scratch. The fine-tuning results for the Recover sub-policy are included in Section A.2 of the supplemental material. We note that Investigate learns different strategies for each adversary, while Recover is relatively agnostic to the attack, and can be directly reused against other red agents. This is understandable, given that the frequency of analyze actions or the placement of decoys is affected by how stealthy or impactful an attacker is, while the recover strategy is attack agnostic, always targeting hosts that present indicators of compromise and performing a remove or restore action.
Last column in Table 2 shows the evaluation results of the master policy that was trained with the fine-tuned sub-policy models. In this case, the master policy performs comparably to (and sometimes even better than) H-MARL Meta that uses sub-policies trained from scratch. For instance, against the Stealthy Red agent, the fine-tuned H-MARL Meta policy attains an average reward of $-162.51$, compared to an average $-184.76$ reward when training from scratch. Our empirical results are supported by previous research, which has shown that pre-trained deep learning models have been proven to generalize better than randomly initialized ones Erhan et al. (2010); Sun et al. (2019).
@@
Figure 6. Investigate sub-policy pre-trained against Aggressive Red is fine-tuned separately against other red agents, performing similarly to the sub-policies trained from scratch against the same red agent.
@@
Table 3. Interpretable metrics for various blue strategies, averaged over 100 evaluation episodes against Default Red.
@@ 5.4. Interpretable Metrics
    • Non-Escalated Hosts: Fraction of hosts with no red root sessions (from total hosts in the network).
@@
    • Mean Time to Recover: Mean number of consecutive time steps that hosts spend in a compromised state.
    • Useful Recoveries: Number of recoveries performed on infected machines (true positives, TP).
    • Wasted Recoveries: Number of recoveries performed on clean machines (false positives, FP).
    • Recovery Error: Fraction of wasted recoveries out of total recoveries: $Err=FP/(TP+FP)$.
    • Recovery Precision: Fraction of useful recoveries out of total recoveries: $RecPrec=TP/(TP+FP)=1-Err$.
@@
    • Red Impact Count: Mean impact count per episode, measuring average number of times the operational service gets compromised.
@@
We evaluate our blue agent strategies according to these metrics in Table 3. We observe a number of insights that are not evident by comparing policies using solely the reward metric. For instance, MARL Decentralized performs more recoveries than other policies, but its recovery precision is only $0.27$. While H-MARL Meta and MARL Decentralized have similar rewards, H-MARL Meta has a much better recovery precision of $0.61$. Moreover, red agents are less successful at impacting the critical services with H-MARL Meta, compared to MARL Decentralized. These indicators show that H-MARL Meta is a more effective defense strategy than the single-policy approach (MARL Decentralized), despite having similar reward. H-MARL Expert has the highest recovery precision across all policies, as its expert master policy selects the Recovery sub-policy when IOCs are present on hosts, a strong indication of host compromise. Our analysis demonstrates the need of using the reward signal in conjunction with additional interpretable metrics that are relevant in the cyber domain.
To better align the reward with the defender’s goals, one can incorporate these new metrics in the reward. As future work, we plan to perform reward shaping to reflect the defender’s objectives.
5.5. What method to use: Expert or Meta?
In our previous experiments, we have seen that H-MARL Expert performs best, as it is guided by a deterministic master policy generated from domain knowledge, specifically: If host IOCs are present, invoke the Recover sub-policy; otherwise Investigate. To show that such expert rules do not generalize in all situations, we consider a scenario with a third Control Traffic sub-policy, in which a defender may block traffic between a clean subnet and another subnet that presents IOCs. However, block actions incur penalties for preventing the remote activity of green users, rendering a deterministic blocking rule unfavorable. We experimented with a probabilistic master policy instead: If host IOCs are present, invoke the Recover sub-policy; otherwise split 75%-25% between Investigate and Control Traffic. With this probabilistic rule, the H-MARL Expert method becomes unstable, while H-MARL Meta performs better, converging fast to a stable performance (Figure 10). These experiments demonstrate the importance of learning the master policy and are discussed in more detail in the supplemental material, Section A.4.
Figure 7. H-MARL with 3 sub-policies: Investigate, Recover, and Control Traffic. Due to the probabilistic expert rule, H-MARL Expert method is unstable, while H-MARL Meta performs well, converging fast to a stable performance.
6. Conclusion
@@
We propose novel hierarchical multi-agent reinforcement learning (MARL) strategies to train multiple blue agents tasked with protecting a network against red agents. Our H-MARL strategy decomposes cyber defense into multiple sub-tasks, trains sub-policies for each sub-task guided by domain expertise, and finally trains a master policy to coordinate sub-policy selection. We evaluate our proposed hierarchical method and compare it against centralized methods in a realistic cyber security environment, CybORG CAGE 4. We demonstrate that our hierarchical policy converges faster than a single PPO policy and achieves higher reward than a hierarchical policy that trains the master policies and sub-policies collectively. We show the generality of our approach across various red agent behavior and introduce a number of interpretable metrics to provide insights to security operators. We leave to future work the question of reward shaping to account for multiple defender objectives.
{acks}
This Research was developed with funding from the Defense
@@
Advanced Research Projects Agency (DARPA), under contract
W912CG23C0031.
Appendix A Supplemental Material
Here we provide some additional experiments and context supplemental to the main body of the paper. In A.1 we provide further details on the CAGE-4 cyber game studied in this paper, in A.2 we show additional transferability experiments on H-MARL Meta, in A.3 we study the potential uses of communication between agents, and in A.4 we further explore the use of Block and Allow Traffic actions in CAGE-4.
A.1. The cyber game
This section presents additional details on the cyber game simulated in CybORG CAGE 4  [^1] https://github.com/cage-challenge/cage-challenge-4/tree/main.
For increased realism, users on enterprise machines are represented by green agents, which are present on every host. Green agents randomly access local and remote services, such as SSHD, MySQL, FTP, etc. Upon compromise, red agents are able to degrade these services, preventing users from completing their work or slowing them down.
@@ A.1. The cyber game
The red team can attack the network through several actions: scan the network to discover new hosts; scan a host to discover active services; exploit a vulnerability to compromise a host; escalate privileges on a compromised node to gain root access; degrade user experience of green agents; impact (stop) the critical Operational technology (OT) service; discover deception (i.e., probe a host to determine if it is running decoy services).
@@
The action sets for green agents (users), blue agents (defenders) and red agents (attackers) are presented in Table 4. Both blue and red agent actions have a specified time duration, ranging from one time step (block/allow traffic) to five time steps for longer actions (restore host). Agents are not allowed to choose another action before the previous one has completed.
Figure 8. The Recover sub-task is rather agnostic to the attack type and can be re-used against other adversaries.
@@
Table 4. Action sets for green agents (users), blue agents (defenders), and red agents (attackers).
@@
A.2. H-MARL Transferability
Figure 8 presents the fine-tuning results of the Recover sub-policy, pre-trained against Aggressive Red, the average-performing attacker. The Recover sub-policy is trained on an observation space consisting of indicators of compromise within a subnet. This sub-policy learns a strong strategy regardless of the attack, and can be directly reused against other red agents.
A.3. Communication
@@
As a case study, we implemented a red agent that chooses external scanning in 50% of the time (once it becomes aware of another subnet), and a blue agent that relies heavily on decoys (90% of blue actions) to detect the adversary during the scanning phase of the attack. Blue agents broadcast 8-bit messages encoding which remote host is accessing their decoys to warn other agents of potential attackers. We uniquely identify the compromised hosts with 3 bits for the subnet number (1-7), and 4 bits for the host index (0-15). Each blue agent decodes the message to check if it refers to hosts from its own subnet. If so, the message provides an indicator of compromise that will be added to the observation vector. Figure 9 shows that there is some benefit of using this method of communication. However, the benefit is small, due to other factors, such as phishing, stealth, and false positives of the Monitor actions.
Figure 9. Blue agents use 8-bit messages to warn other team members of potential compromised hosts. This communication strategy shows some benefit over the case when no communication is used. (MARL Decentralized training)
Figure 10. H-MARL with 3 sub-policies: Investigate, Recover, and Control Traffic. Due to the probabilistic expert rule, H-MARL Expert method is unstable, while H-MARL Meta performs well, converging fast to a stable performance.
Figure 11. H-MARL with 3 sub-policies: Investigate, Recover, and Control Traffic, after removing the penalty on failed user access (no degrade). All sub-policies are fine-tuned in this new context, against a new red that attempts frequent remote exploits into other subnets. The H-MARL Expert has regained a stable high performance and H-MARL Meta has similar performance.
A.4. Traffic Control
@@
For our current experiments, we use the following expert rule to train the sub-policies (Algorithm 1), and leave further research into other expert rules for future work: If indicators of compromise are present on hosts, call the Recover sub-policy; otherwise, randomly choose Investigate for 75% of the time, and Control Traffic for the remaining 25%. Thus, the Investigate sub-policy is assigned more weight, as we expect it to be useful more frequently. In fact, for best performance, we use (keep fixed) the Investigate and Recover sub-policies trained previously (see paper), since they have been learned with well-defined expert knowledge, and only train the Control Traffic sub-policy, using the rule specified above.
Next, we follow Algorithm 2 to train the Master policy using the three pre-trained sub-policies. Figure 10 shows the reward as training progresses during Algorithm 1 (using the expert rule), and Algorithm 2 (train the master). We observe the instability of the expert, which has been trained with a probability-based rule, reinforcing the importance of learning the master policy. The H-MARL Meta algorithm is more stable, as the master policy learns how to combine the sub-policies to solve the meta-task, and is not restricted by a fixed, deterministic rule.
Turning off Degrade Service. In our next set of experiments illustrated in Figure 11, we fine-tuned the sub-policies after turning off the reward penalty of green agents being affected by failed service access. We also used the modified red agent introduced in Section A.3 of the supplemental material, which performs remote scanning into other subnets more often, and can still collect rewards by impacting the critical OT service. With block actions being now useful at preventing red agents from spreading, without incurring penalties, H-MARL Expert regains a stable, high performance, as expected. H-MARL Meta achieves similar performance with H-MARL Expert in this setting. The blue agents are using 4×4$\times$4 × more block actions when the failed user access penalty has been removed.

### v3

@@ 5.3 H-MARL Transferability
In this section, we explore the possibility of knowledge transfer (Weinshall2018, ; pmlr-v80-wei18a, ; zhuang2020transferlearning, ; NekoeiBCC21, ) – given a defense strategy that has already been trained, can we use it to accelerate the learning of a new defense that protects against a different adversary? The adversaries in this study are stationary randomized agents, which consistently apply the same probabilistic rules to make decisions. We leave as future work the study of H-MARL transferability to evolving adversaries that adapt their policies based on opponent behavior or environment feedback. We adapt the pre-trained sub-policy models to new attacks via fine-tuning, a well-established powerful transfer method for deep models (zhuang2020transferlearning, ; Sun2019, ) that is significantly less costly than training from scratch. In our current design, we only fine-tune the sub-policies for a few iterations. Next, we fully train a new master policy, using the tuned sub-policies.
@@
Figure 6: The Investigate sub-policy pre-trained against Aggressive Red is fine-tuned separately against other red agents, performing similarly to training from scratch.
Figure 6 presents the fine-tuning results of the Investigate sub-policy, pre-trained against Aggressive Red, the average-performing attacker. A short fine-tuning is enough to adapt to a new adversary, resulting in learning curves similar to those obtained when training from scratch. The fine-tuning results for the Recover sub-policy are included in Section S-5 the Supplemental Material. We note that Investigate learns different strategies for each adversary (due to different patterns of suspicious events), while Recover is relatively agnostic to the attack (due to working on clear indicators of compromise) and can be directly reused against other red agents.
Last row in Table 2 shows the evaluation results of the master policy that was trained with the fine-tuned sub-policy models. In this case, the master policy performs comparably to (and sometimes even better than) H-MARL Meta that uses sub-policies trained from scratch. For instance, against the Stealthy Red agent, the fine-tuned H-MARL Meta policy attains an average reward of 162.51-162.51, compared to an average 184.76-184.76 reward when training from scratch. Our empirical results are supported by previous research, which has shown that pre-trained deep learning models have been proven to generalize better than randomly initialized ones (Erhan2010, ; Sun2019, ).
5.4 What method to use: Expert or Meta?
In our previous experiments, we have seen that H-MARL Expert performs best, as it is guided by a deterministic master policy generated from domain knowledge, specifically: If host IOCs are present, invoke the Recover sub-policy; otherwise Investigate. To show that such expert rules do not generalize in all situations, we consider a scenario with a third sub-policy, Control Traffic, in which a defender may choose to block traffic between subnets.
However, block actions incur penalties for preventing the remote activity of green users, rendering a deterministic blocking rule unfavorable. We instead experimented with a probabilistic master policy: If host IOCs are present, invoke the Recover sub-policy; otherwise split 75%-25% between Investigate and Control Traffic. With this probabilistic rule, the H-MARL Expert method becomes unstable, while H-MARL Meta performs better and converges quickly to a stable performance (see Figure 7). This demonstrates the importance of learning the master policy and is discussed in more detail in Section S-4 of the Supplemental Material.
Figure 7: H-MARL with 3 sub-policies: Investigate, Recover, and Control Traffic. Due to the probabilistic expert rule, H-MARL Expert method is unstable, while H-MARL Meta performs well, converging fast to a stable performance.
@@ 5.4 What method to use: Expert or Meta?
Table 3: Interpretable metrics for various blue strategies against Default Red, 100-episode averages.
@@ 5.5 Interpretable Metrics
    • Non-Escalated Hosts: Fraction of hosts with no red root sessions.
@@
    • Mean Time to Recover: Mean number of consecutive steps spent in a compromised state.
    • Useful Recoveries: Recoveries performed on infected machines (true positives, TP).
    • Wasted Recoveries: Recoveries performed on clean machines (false positives, FP).
    • Recovery Error: Err = FP / (TP + FP).
    • Recovery Precision: TP / (TP + FP) = 1-Err.
@@
    • Red Impact Count: Number of times the OT service is impacted becoming unavailable.
@@
We evaluate our blue agent strategies according to these metrics in Table 3. We observe a number of insights that are not evident by comparing policies using solely the reward metric. For instance, MARL Decentralized performs more recoveries than other policies, but its recovery precision is only $0.27$. While H-MARL Meta and MARL Decentralized have similar rewards, H-MARL Meta has a much better recovery precision of $0.61$. Moreover, red agents are less successful at impacting the critical services with H-MARL Meta, compared to MARL Decentralized. These indicators show that H-MARL Meta is a more effective defense strategy than the single-policy approach (MARL Decentralized), despite having similar reward. H-MARL Expert has the highest recovery precision across all policies, as its expert master policy selects the Recovery sub-policy when IOCs are present on hosts, a strong indication of host compromise. Our analysis demonstrates the need of using the reward signal in conjunction with other metrics that are relevant in the cyber domain. To better align the reward with the defender’s goals, one can also incorporate these new metrics in the reward.
6 Conclusion
@@
We propose novel hierarchical multi-agent reinforcement learning (MARL) strategies to train multiple blue agents tasked with protecting a network against red agents. Our H-MARL strategy decomposes cyber defense into multiple sub-tasks, trains sub-policies for each sub-task guided by domain expertise, and finally trains a master policy to coordinate sub-policy selection. We evaluated our proposed hierarchical methods (Expert and Meta) and compared them against standard decentralized and centralized MARL in a realistic cyber security environment, CybORG CAGE 4. We demonstrated that our hierarchical method converges faster than a single PPO policy and generalizes across various red agent behavior, while H-MARL Expert consistently performs better than the baselines.
Acknowledgments
@@
This research was funded by the Defense Advanced Research Projects Agency (DARPA), under contract W912CG23C0031.
Supplementary Materials
S-1 The Cyber Game
@@
This section presents additional details on the cyber game simulated in CybORG CAGE 4  [^2] https://github.com/cage-challenge/cage-challenge-4/tree/main. For increased realism, users on enterprise machines are represented by green agents, which are present on every host. Green agents randomly access local and remote services, such as SSHD, MySQL, FTP, etc. Upon compromise, red agents are able to degrade these services, preventing users from completing their work or slowing them down. The red team can attack the network through several actions: scan the network to discover new hosts; scan a host to discover active services; exploit a vulnerability to compromise a host; escalate privileges on a compromised node to gain root access; degrade user experience of green agents; impact (stop) the critical Operational technology (OT) service; discover deception (i.e., probe a host to determine if it is running decoy services).
@@
The action sets for green agents (users), blue agents (defenders) and red agents (attackers) are presented in Table S-1. Both blue and red agent actions have a specified time duration, ranging from one time step (block/allow traffic) to five time steps for longer actions (restore host). Agents are not allowed to choose another action before the previous action had completed.
@@
Table S-1: Action sets for green agents (users), blue agents (defenders), and red agents (attackers).
@@
The red agent in CAGE 4 is rather sophisticated and randomized, not deterministic, and is modeled by:
  • State transition diagram: The adversary follows a state transition diagram with eight states and nine transitions (that is, actions) that defines the allowed subset of actions based on the state of the host.
  • State transition probability matrix: The adversary’s actions are randomized, using a state transition probability matrix to choose among the subset of possible actions in each state. For example, in State S (i.e., a Service was discovered on a host), the Red agent can choose to Discover Remote Systems with 0.25 probability, Discover Deception with 0.25 probability, or Exploit Remote Service with 0.5 probability.
We vary the probabilities in the state transition probability matrix to create additional red agent variants and cover a wider range of Red behaviors (fully Aggressive, Stealthy, or Impact). These new attack vectors explore how two important characteristics of any cyber attack, namely the speed/stealth of discovering vulnerabilities, and the attacker’s objective affect the attack success.
S-2 Communication and Cooperation
@@
Figure S-1: Blue agents use 8-bit messages to warn other team members of potential compromised hosts. This communication strategy shows some benefit over the case when no communication is used. (MARL Decentralized training)
As a case study, we implemented a red agent that chooses external scanning in 50% of the time (once it becomes aware of another subnet), and a blue agent that relies heavily on decoys (90% of blue actions) to detect the adversary during the scanning phase of the attack. Blue agents broadcast 8-bit messages encoding which remote host is accessing their decoys to warn other agents of potential attackers. We uniquely identify the compromised hosts with 3 bits for the subnet number (1-7), and 4 bits for the host index (0-15). Each blue agent decodes the message to check if it refers to hosts from its own subnet. If so, the message provides an indicator of compromise that will be added to the observation vector. Figure S-1 shows that there is some benefit of using this method of communication. However, the benefit is small, due to other factors, such as phishing, stealth, and false positives of the Monitor actions.
S-3 H-MARL Pipeline
Meta actions describe a class of actions (e.g., Restore, Investigate, Control Traffic), which the master policy selects. This partitioning abstracts away additional details, such as what machine to restore, making training and generalization easier for the master policy. Sub-policies describe the policies that choose the primitive action ultimately executed in the environment (e.g., restore host 13). In effect the master policy chooses a meta action, and then samples a primitive action from the respective sub-policy.
Figure S-2: H-MARL Expert Pipeline. The Master uses expert rules to choose a sub-policy that steps through the environment. Sub-policy training is guided by the reward from the environment.
An overview of the H-MARL Expert pipeline is shown in Figure S-2. Upon receiving an observation the expert master uses pre-defined rules to choose a meta-action indexed by $i$ The observation is then processed into the respective observation space of sub-policy $i$, which chooses the final primitive action to step through the environment. This approach aims to learn sub-policies that are specialized for a single task.
Figure S-3: H-MARL Meta Pipeline. The Master learns a probability distribution over meta-actions. The Master training uses frozen sub-policies and is guided by the reward from the environment.
An overview of the H-MARL Meta pipeline is shown in Figure S-3. The meta master policy learns to reason on a higher-level about the decisions it can make by using state abstractions (e.g. are IOCs present in the network?). During the training of the meta master, the sub-policies are kept frozen. The master learns a probability distribution over the meta-actions, guided by the reward from the environment.
S-4 Traffic Control
@@
Figure S-4: H-MARL with 3 sub-policies: Investigate, Recover, and Control Traffic. Due to the probabilistic expert rule, H-MARL Expert method is unstable, while H-MARL Meta performs well, converging fast to a stable performance.
Figure S-5: H-MARL with 3 sub-policies: Investigate, Recover, and Control Traffic, after removing the penalty on failed user access (no degrade). All sub-policies are fine-tuned in this new context, against a new red that attempts frequent remote exploits into other subnets. The H-MARL Expert has regained a stable high performance and H-MARL Meta has similar performance.
@@
For our current experiments, we use the following expert master policy to train the sub-policies (Algorithm 1), and leave further research into other expert rules for future work: If indicators of compromise are present on hosts, call the Recover sub-policy; otherwise, randomly choose Investigate for 75% of the time, and Control Traffic for the remaining 25%. Thus, the Investigate sub-policy is assigned more weight, as we expect it to be useful more frequently. In fact, for best performance, we use (keep fixed) the Investigate and Recover sub-policies trained previously (see paper), since they have been learned with well-defined expert knowledge, and only train the Control Traffic sub-policy, using the rule specified above.
Next, we follow Algorithm 2 to train the Master policy using the three pre-trained sub-policies. Figure S-4 shows the reward as training progresses during Algorithm 1 (using the expert rule), and Algorithm 2 (train the master). We observe the instability of the expert, which has been trained with a probability-based rule, reinforcing the importance of learning the master policy. The H-MARL Meta algorithm is more stable, as the master policy learns how to combine the sub-policies to solve the meta-task, and is not restricted by a fixed, deterministic rule.
Turning off Degrade Service. In our next set of experiments illustrated in Figure S-5, we fine-tuned the sub-policies after turning off the reward penalty of green agents being affected by failed service access. We also used the modified red agent introduced in Section S-2 of the supplemental material, which performs remote scanning into other subnets more often, and can still collect rewards by impacting the critical OT service. With block actions being now useful at preventing red agents from spreading, without incurring penalties, H-MARL Expert regains a stable, high performance, as expected. H-MARL Meta achieves similar performance with H-MARL Expert in this setting. The blue agents are using $4\times$ more block actions when the failed user access penalty has been removed.
S-5 H-MARL Transferability
Figure S-6 presents the fine-tuning results of the Recover sub-policy, pre-trained against Aggressive Red, the average-performing attacker. The Recover sub-policy is trained on an observation space consisting of indicators of compromise within a subnet. This sub-policy learns a strong strategy regardless of the attack, and can be directly reused against other red agents.
Figure S-6: The Recover sub-task is rather agnostic to the attack type and can be re-used against other adversaries.